# Are You Doubtful? Oh, It Might Be Difficult Then! Exploring the Use of Model Uncertainty for Question Difficulty Estimation

### v1

@@ Abstract
In an educational setting, an estimate of the difficulty of multiple-choice questions (MCQs), a commonly used strategy to assess learning progress, constitutes very useful information for both teachers and students. Since human assessment is costly from multiple points of view, automatic approaches to MCQ item difficulty estimation are investigated, yielding however mixed success until now. Our approach to this problem takes a different angle from previous work: asking various Large Language Models to tackle the questions included in two different MCQ datasets, we leverage model uncertainty to estimate item difficulty. By using both model uncertainty features as well as textual features in a Random Forest regressor, we show that uncertainty features contribute substantially to difficulty prediction, where difficulty is inversely proportional to the number of students who can correctly answer a question. In addition to showing the value of our approach, we also observe that our model achieves state-of-the-art results on the BEA publicly available dataset.
@@
tcb@breakable
Are You Doubtful? Oh, It Might Be Difficult Then!
 Exploring the Use of Model Uncertainty for Question Difficulty Estimation
Leonidas Zotos University of Groningen l.zotos@rug.nl                        Hedderik van Rijn University of Groningen d.h.van.rijn@rug.nl                        Malvina Nissim University of Groningen m.nissim@rug.nl
1 Introduction
Multiple-Choice Questions (MCQs) are commonly used as a form of assessment across educational levels. This is not surprising, as they are trivial to grade and can effectively assess a student’s knowledge, as long as they are designed well Gierl et al. (2017). Naturally, an aspect that significantly affects an MCQ’s quality is its difficulty. Intuitively, items that are too easy do not sufficiently challenge students, while very difficult items lead to frustration and demotivation Papoušek et al. (2016) impairing the learning process. However, estimating an item’s difficulty is not trivial. In fact, students, and especially teachers, are not great at estimating how many of the test-takers will select the correct answer, given a question van de Watering and van der Rijt (2006). While field-testing question items is a viable solution, it is usually expensive, both in terms of time and resources.
@@ 1 Introduction
Computational methods, including Large Language Models (LLMs), have had some success in assessing the difficulty of MCQs AlKhuzaey et al. (2024). At the same time, the task remains challenging, as shown by a recent shared task on automated difficulty prediction for MCQs Yaneva et al. (2024), where most submitted systems performed barely above some simple baselines. The goal of the current work is to tackle the task of item difficulty estimation using a minimal experimental setup showcasing the usefulness of model uncertainty for this task. We do this by obtaining a score for the uncertainty LLMs exhibit when answering a variety of MCQs and use it, in combination with basic text and semantic features, to train an interpretable regressor model. This expands on previous findings which showed a correlation between model and student perceived difficulty (Zotos et al., 2024), paired with the intuition that both syntactic and semantic features are integral to this task AlKhuzaey et al. (2024). We focus on factual MCQs, as this type of assessment is less subjective compared to open-ended questions, while still offering more complexity than simple True/False questions [^1] In True/False questions, a statement needs to be assessed as correct or incorrect, with a random chance of 50%. In contrast, MCQs can follow various formulations and the distractors play a significant role.. Lastly, this choice is also motivated by dataset availability, as explained in Section 3.
It is worthwhile explicitly mentioning that in the current work, the term "uncertainty" is used to encompass both 1st token probability and choice-order probability metrics (see Section 4.2 for details.) These measures are taken to broadly represent the inverse of model confidence. While accurately determining the uncertainty of an LLM is an open field of research, previous research suggests that both 1st token probability Plaut et al. (2024) and choice order probability Zotos et al. (2024) correlate well with model correctness in the MCQs setup. These findings also hold in the current experimental setup, as shown in Appendix A.
@@
The contribution of our work is twofold. First, thanks to extensive experiments with a variety of LLMs and feature analysis using an interpretable model (Random forest Regressor), we showcase that model uncertainty is a useful proxy for item difficulty estimation on two different question sets assessing factual knowledge. Second, as a byproduct of our experiments investigating model uncertainty we yield a model which achieves best results to date on the BEA 2024 Shared Task dataset. This model, together with all experimental code, is made available to the community for replicability and future extensions. We believe that our conceptual insight (model uncertainty as a useful signal for item difficulty), as well as our practical contribution in terms of an existing modular system, will foster further improvements in the task of MCQ automatic difficulty estimation, which is core in the educational setting. [^2] Code will become available upon acceptance.
@@ 2 Related work
The task of estimating the difficulty of MCQ items has been explored from various viewpoints in the literature AlKhuzaey et al. (2024). Most commonly this task is tackled by training a model on a set of syntactic (Perkins et al., 1995; Ha et al., 2019, e.g.,) and/or semantic features (Xue et al., 2020; Hsu et al., 2018, e.g.,). Furthermore, the majority of studies focus on the field of Language learning (Bi et al., 2021; He et al., 2021, e.g.,) which is inherently different to factual knowledge examinations. While the task of difficulty estimation has been widely explored, it remains challenging as was also seen in the recent "Building Educational Applications" (BEA) shared task on "Automated Prediction of Item Difficulty and Item Response Time", where simple baselines were overall only marginally beaten Yaneva et al. (2024). In this task, a variety of approaches were explored with the focus ranging from architectural changes to data augmentation techniques. Notably, the best performing team (EduTec) used a combination of model optimisation techniques, namely scalar mixing, rational activation and multi-task learning (leveraging the provided response time measurements also provided in the BEA dataset) Gombert et al. (2024).
@@
Most similar to our work is the study by Loginova et al. (2021), who also explore the use of confidence of language models to estimate question difficulty. While similarities exist, the current research deviates considerably from this study. Their focus is on language comprehension, which differs from our emphasis on factual knowledge assessment. Furthermore, while their focus is on Encoder-Only models, ours is on Decoder-Only models, which incorporate greater amounts of factual knowledge as a byproduct of their language modeling objective Zhao et al. (2023).
@@
More broadly, there is an emerging "LLM-as-a-judge" field of research, which, in general terms, explores the possibility of using powerful LLMs as a substitute for human annotation Zheng et al. (2023); Pan et al. (2024). For the task of question difficulty estimation, this paradigm has been explored in the context of language comprehension by Raina and Gales (2024) with some success. However, our preliminary research suggests that this approach does not perform well for the two datasets used in the current study (see Appendix B), thereby motivating further exploration in novel approaches.
@@
The present work directly builds on the work by Zotos et al. (2024), where a variety of analyses showed promising results on correlating human and machine perceived difficulty. We take this one step further, by testing a battery of different LLMs on item difficulty estimation using their uncertainty as a signal, focusing on two distinct question sets assessing factual knowledge.
3 Data
The two factual knowledge MCQ datasets that we use in our experiments are described more in detail in the following subsections. The first is a dataset on the domain of Biopsychology that is not publicly available, while the second is the publicly available dataset used in the BEA 2024 Shared Task Yaneva et al. (2024). For brevity, we refer to the "Biopsychology" and "BEA" datasets respectively. Our choice is driven by the requirement of having question-sets along with students selection rates (serving as proxies for item difficulty scores). Considering that, to the best of our knowledge, the BEA dataset is the only publicly available resource satisfying this requirement, we also use a non-publicly available dataset. This choice is in line with the observation by AlKhuzaey et al. (2024), who note that most studies tackling this task resort to using private datasets.
We use first the Biopsychology dataset to extend the experiments in (Zotos et al., 2024), and then the BEA dataset to evaluate whether our approach generalises to a different dataset (for which many other systems exist and can be compared to). As will be explained in Sections 3.1 and 3.2, even though both question sets assess factual knowledge, they also vary in multiple aspects, for example question formulation, number of distractors and knowledge specificity. Furthermore, to facilitate comparison with the findings from the BEA 2024 Shared Task, we use the train/test split as provided in the shared task itself (70% training and 30% test samples). The same proportions are also used for the Biopsychology dataset, as shown in Table 1.
The item difficulty labels differ between the two datasets. In the Biopsychology dataset, difficulty is measured by the proportion of students who answered correctly (a higher value indicates an easier question). In contrast, the BEA dataset originally uses the inverse difficulty measurement, where a higher difficulty label signifies that fewer students answered correctly. Additionally, a linear transformation is also applied on the target labels of the BEA dataset. While this difference does not affect our approach, as in both cases difficulty is conceptually expressed by cumulative student performance, to allow easier interpretation of our results we have transformed the BEA difficulty scores to their complements such that they also reflect the proportion of correct responses per question.
Dataset Train Test Total
Biopsychology 573 246 819
BEA 466 201 667
Table 1: Train and Test splits as used in our experiments. For BEA, we use the splits as provided in the competition Yaneva et al. (2024). For Biopsychology, we randomly sampled the questions, keeping the same percentage of training/testing samples as in BEA.
3.1 Biopsychology
[…]
3.2 BEA 2024
The BEA question set was used in the United States Medical Licensing Examination (USMLE®) and was developed by the National Board of Medical Examiners (NBME®) and Federation of State Medical Boards (FSMB) Yaneva et al. (2024). It consists of $667667$ MCQs, each answered by more than $300300$ medical school students. In contrast to the Biopsychology dataset, the questions follow strict guidelines (e.g., fixed structure, absence of misleading or redundant information in the question) and are presented with up to nine distractor choices, with the majority of the questions having five ($525525$ items) or six distractors ($7171$ items). An example instance is provided in Table 2. As can be seen, questions of this dataset are significantly longer ($755755$ characters compared to $103103$ characters for the Biopsychology set) and are of technical nature.
@@ 3.2 BEA 2024
BEA A 65-year-old woman comes to the physician for a follow-up examination after blood pressure measurements were 175/105 mm Hg and 185/110 mm Hg 1 and 3 weeks ago, respectively. She has well-controlled type 2 diabetes mellitus. Her blood pressure now is 175/110 mm Hg. Physical examination shows no other abnormalities. Antihypertensive therapy is started, but her blood pressure remains elevated at her next visit 3 weeks later. Laboratory studies show increased plasma renin activity; the erythrocyte sedimentation rate and serum electrolytes are within the reference ranges. Angiography shows a high-grade stenosis of the proximal right renal artery; the left renal artery appears normal. Which of the following is the most likely diagnosis? Atherosclerosis
@@
Table 2: Examples questions from the Biopsychology and BEA datasets. Correct answer in green.
@@ 4 Approach
Given an MCQ, the task is to predict the proportion of students that select the correct choice [^3] This is also known as the p-value van de Watering and van der Rijt (2006).. An MCQ item consists of the stem/question, a single correct choice/answer and a number of incorrect choices/distractors (also known as “foils").
Figure 1 illustrates our approach to this task. Our design is centered around a simple Random Forest Regressor [^4] As provided by the Scikit-Learn Library, using the default hyper-parameters Pedregosa et al. (2011). which receives as input a vectorised representation of the MCQ, as well the uncertainty of multiple LLMs answering the same MCQ [^5] Simple vector concatenation is used to combine the text and uncertainty features.. We opted for a relatively simple Random Forest Regressor due to its interpretability compared to more complex architectures, while still effectively demonstrating the usefulness of model uncertainty in this context. As features, we use Textual Features and Model Uncertainty, as described in the following sections.
@@ 4.1 Textual Features
TF-IDF Scores capture how important a word is to a document within a collection by balancing its frequency in that specific document against its rarity across all documents Sparck Jones (1972). In the current context, we consider each question item (along with its choices) as a single document. To capture multi-word technical terms, such as "interstitial fibrosis", our analysis considers both individual words (unigrams) and two-word combinations (bigrams). Furthermore, we disregard terms that appear in more than 75% of documents, and only use the $10001000$ most important features (as determined by the TF-IDF values) to increase efficiency.
@@
Word embeddings are a technique whereby words are encoded as dense vectors in a continuous vector space, capturing semantic relationships between words. We evaluate two embedding approaches: General BERT embeddings Devlin et al. (2019) and domain-specific Bio-Clinical BERT embeddings Alsentzer et al. (2019), relevant to the topics addressed in the MCQ datasets we use. The Bio-Clinical BERT embeddings, previously also employed by team ITEC in the BEA 2024 shared task Tack et al. (2024), offer specialized medical domain text encoding that potentially encapsulates more accurately the semantic content of each question item. Both techniques yield a 768-dimensional vector representation.

### v2

@@ Abstract
In an educational setting, an estimate of the difficulty of multiple-choice questions (MCQs), a commonly used strategy to assess learning progress, constitutes very useful information for both teachers and students. Since human assessment is costly from multiple points of view, automatic approaches to MCQ item difficulty estimation are investigated, yielding however mixed success until now. Our approach to this problem takes a different angle from previous work: asking various Large Language Models to tackle the questions included in three different MCQ datasets, we leverage model uncertainty to estimate item difficulty. By using both model uncertainty features as well as textual features in a Random Forest regressor, we show that uncertainty features contribute substantially to difficulty prediction, where difficulty is inversely proportional to the number of students who can correctly answer a question. In addition to showing the value of our approach, we also observe that our model achieves state-of-the-art results on the USMLE and CMCQRD publicly available datasets.
keywords:
item difficulty estimation, model uncertainty, multiple choice questions
1 Introduction
Multiple-Choice Questions (MCQs) are commonly used as a form of assessment across educational levels. This is not surprising, as they are trivial to grade and can effectively assess a student’s knowledge, as long as they are designed well [6]. Naturally, an aspect that significantly affects an MCQ’s quality is its difficulty (broadly determined by the students’ success answering the question item). Intuitively, items that are too easy do not sufficiently challenge students, while very difficult items lead to frustration and demotivation impairing the learning process [17]. However, estimating an item’s difficulty is not trivial. In fact, students, and especially teachers, are not great at estimating how many of the test-takers will select the correct answer, given a question [30]. While field-testing question items is a viable solution, it is usually expensive, both in terms of time and resources.
@@ 1 Introduction
Approach overview: Predicting difficulty of Multiple-Choice Question items using textual features and uncertainty of LLM test-takers. Refer to caption
@@
Computational methods, including Large Language Models (LLMs), have had some success in assessing the difficulty of MCQs [1]. At the same time, the task remains challenging, as shown by a recent shared task on automated difficulty prediction for MCQs [36], where most submitted systems performed barely above some simple baselines. The goal of the current work is to tackle the task of item difficulty estimation using a minimal experimental setup showcasing the usefulness of model uncertainty for this task. We do this by obtaining a score for the uncertainty LLMs exhibit when answering a variety of MCQs and use it, in combination with basic text and semantic features, to train a regressor model. This expands on previous findings which showed a correlation between model and student perceived difficulty [40], paired with the intuition that both syntactic and semantic features are integral to this task [1]. We focus on factual MCQs, as they provide more objective assessment than open-ended questions, while still offering more complexity than True/False questions where the a baseline random chance is 50%. In contrast, MCQs can follow various formulations and the incorrect choices play a significant role. Lastly, this choice is also motivated by dataset availability, as explained in Section 3.
It is worthwhile explicitly mentioning that in the current work, the term “uncertainty" is used to encompass both 1st token probability and choice-order probability metrics (see Section 4.2 for details.) These measures are taken to broadly represent the inverse of model confidence. While accurately determining the uncertainty of an LLM is an open field of research, previous research suggests that both 1st token probability [22] and choice order probability [40] correlate well with model correctness in the MCQ setup. These findings also hold in the current experimental setup, as shown in Appendix A.
@@
The contribution of our work is twofold. First, thanks to extensive experiments with a variety of LLMs and feature analysis using a regressor model (Random Forest Regressor), we showcase that model uncertainty is a useful proxy for item difficulty estimation on three different question sets assessing both factual knowledge and reading comprehension. Second, as a byproduct of our experiments investigating model uncertainty we yield a model which achieves best results to date on the BEA 2024 Shared Task dataset as well as the CMCQRD dataset. This model, together with all experimental code, is made available to the community for replicability and future extensions. We believe that our conceptual insight (model uncertainty as a useful signal for item difficulty), as well as our practical contribution in terms of an existing modular system, will foster further improvements in the task of MCQ automatic difficulty estimation, which is core in the educational setting. [^1] Code available at: https://github.com/LeonidasZotos/Are-You-Doubtful-Oh-It-Might-Be-Difficult-Then
@@ 2 Related work
The task of estimating the difficulty of MCQ items has been explored from various viewpoints in the literature [1]. Most commonly this task is tackled by training a model on a set of syntactic [20, 8] and/or semantic features [34, 10]. Furthermore, the majority of studies focus on the field of Language learning [3, 9] which is inherently different to factual knowledge or reading comprehension examinations. While the task of difficulty estimation has been widely explored, it remains challenging as was also seen in the recent “Building Educational Applications" (BEA) shared task on “Automated Prediction of Item Difficulty and Item Response Time", where simple baselines were overall only marginally beaten [36]. In this task, a variety of approaches were explored with the focus ranging from architectural changes to data augmentation techniques. Notably, the best performing team (EduTec) used a combination of model optimisation techniques, namely scalar mixing, rational activation and multi-task learning (leveraging the provided response time measurements also provided in the USMLE dataset) [7].
@@
Most similar to our work is the study by Loginova et al. [12], who also explore the use of confidence of language models to estimate question difficulty. While similarities exist, the current research deviates considerably from this study. Whereas Loginova et al. focus on training and calibrating Encoder-Only Question-Answering models, we instead examine “off-the-shelf" Decoder-Only models, which inherently incorporate a greater amount of factual knowledge as a byproduct of their language modeling objective [37]. Additionally, we broaden the research scope beyond language comprehension to include factual knowledge understanding, through the use of three datasets that assess knowledge across different education levels. Finally, rather than relying on proxied comparative assessments – where assessors classify questions as “easy" or “difficult" – we leverage fine-grained, continuous difficulty labels, such as the proportion of students answering a question correctly.
@@
More broadly, there is an emerging “LLM-as-a-judge" field of research, which, in general terms, explores the possibility of using powerful LLMs as a substitute for human annotation [39, 16]. For the task of question difficulty estimation, this paradigm has been explored in the context of language comprehension by with some success [23]. More recently Raina et al. achieved good results in the USMLE and CMCQRD datasets through comparative assessments (i.e., given two question items, the LLM’s task is to determine which is more difficult) [24].
@@
The present work is also influenced by the work by Zotos et al., where a variety of analyses showed a weak, but statistically significant, correlation between human and machine perceived difficulty [40]. We take this one step further, by testing a battery of different LLMs on item difficulty estimation using their uncertainty as a signal, focusing on three distinct question sets assessing both factual knowledge and reading comprehension.
Table 1: Examples questions from the Biopsychology, USMLE and CMCQRD datasets. Correct answer in green. Two examples are given from the Biopsychology dataset to illustrate the phrasing variability.
@@
USMLE A 65-year-old woman comes to the physician for a follow-up examination after blood pressure measurements were 175/105 mm Hg and 185/110 mm Hg 1 and 3 weeks ago, respectively. She has well-controlled type 2 diabetes mellitus. Her blood pressure now is 175/110 mm Hg. Physical examination shows no other abnormalities. Antihypertensive therapy is started, but her blood pressure remains elevated at her next visit 3 weeks later. Laboratory studies show increased plasma renin activity; the erythrocyte sedimentation rate and serum electrolytes are within the reference ranges. Angiography shows a high-grade stenosis of the proximal right renal artery; the left renal artery appears normal. Which of the following is the most likely diagnosis? Atherosclerosis
@@
CMCQRD Mum caught Jess by the arm. ‘Come with me,’ she said. Jess followed her through to the study and there on the table, propped against the wall, was an unframed painting, unmistakably one of Grandpa’s, yet unlike anything he had done before, and clearly nowhere near finished. ‘Do you know anything about this?’ said Mum. Jess shook her head. ‘I’ve never seen it before. I didn’t know he was working on anything.’ […] And it was hard to imagine anyone, even a goddess, having any influence over someone as wilful as Grandpa. ‘He doesn’t need me to inspire him,’ she said. ‘He’s been painting all his life.’
What impression is given of Jess’s grandfather in the final paragraph?
ⓐ He lacks confidence in his ability.
ⓑ He values Jess’s opinions.
ⓒ He has a strong character.
ⓓ He finds it hard to concentrate.
3 Data
The three MCQ datasets that we use in our experiments are described more in detail in the following subsections. The first is a dataset on the domain of Biopsychology that is not publicly available. The second is the publicly available dataset used in the BEA 2024 Shared Task [36]. Lastly, we also use the “Cambridge Multiple-Choice Questions Reading Dataset" [15]. For brevity, we refer to the “Biopsychology", “USMLE" and “CMCQRD" datasets respectively. Our choice is driven by the requirement of having question-sets along with students selection rates (serving as proxies for item difficulty scores). Considering that, to the best of our knowledge, the USMLE and CMCQRD datasets are the only publicly available resources satisfying this requirement. Additionally, we use a non-publicly available dataset as a complement. This choice is in line with the observation by AlKhuzaey et al., who note that most studies tackling this task resort to using private datasets [1].
Table 2: Train and Test splits as used in our experiments. For USMLE, we use the splits as provided in the competition [36]. For the Biopsychology and CMCQRD datasets, we randomly sampled the questions, keeping the same percentage of training/testing samples as in USMLE. For the CMCQRD, we also ensured that text passages present in the train set were not present in the test set.
Dataset Train Test Total Public?
Biopsychology 573 246 819
USMLE 466 201 667
CMCQRD 552 241 793
As will be explained in Sections 3.1 through 3.3, the three datasets vary in multiple aspects, for example question formulation, number of incorrect choices (also known as distractors) and knowledge specificity. Furthermore, to facilitate comparison with the findings from the BEA 2024 Shared Task, we use the train/test split as provided in the shared task itself (70% training and 30% test samples). The same proportions are also used for the Biopsychology and CMCQRD datasets, as shown in Table 2.
The item difficulty labels differ between the three datasets. In the Biopsychology and CMCQRD datasets, difficulty is measured by the proportion of students who answered correctly (a higher value indicates an easier question). This measure is commonly known as the p-value of a question item [30]. In contrast, the USMLE dataset originally uses the inverse difficulty measurement, where a higher difficulty label signifies that fewer students answered correctly. Additionally, a linear transformation is also applied on the target labels of the USMLE dataset. While this difference does not affect our approach, as in both cases difficulty is conceptually expressed by cumulative student performance, to allow easier interpretation of our results we have transformed the USMLE difficulty scores to their complements such that they also reflect the proportion of correct responses per question.
[…]
Biopsychology dataset is least balanced, with most questions being very easy. USMLE has fairly spread out p-values. Refer to caption
Figure 2: Distribution of question difficulty, based on the proportion of students answering each question correctly. For the USMLE dataset, only transformed p-values are available.
Finally, CMCQRD is also equipped with an Item Response Theory (IRT)-based metric, specifically scaled single-parameter Rasch Model difficulty estimates. To the best of our knowledge, while previous results exist using this IRT-based metric [24], no prior work has attempted to estimate p-values for this dataset. In this work, we conduct experiments independently estimating both the p-value and the IRT-based metric, achieving state-of-the-art results with the latter, where prior results are also available.
Figure 2 illustrates the distribution of difficulty across the three datasets (focusing on the p-value for the CMCQRD and not the IRT metric). As shown, while the datasets contain questions of varying difficulty levels, they are generally skewed toward easier questions. This trend is particularly evident in the Biopsychology dataset, where 81% of questions were answered correctly by at least 60% of the student population.
3.1 Biopsychology
[…]
3.2 USMLE
The United States Medical Licensing Examination (USMLE) question set was developed by the National Board of Medical Examiners (NBME) and Federation of State Medical Boards (FSMB) [36]. It consists of $667667$ MCQs, each answered by more than $300300$ medical school students. In contrast to the Biopsychology dataset, the questions follow strict guidelines (e.g., fixed structure, absence of misleading or redundant information in the question) and are presented with up to nine distractor choices, with the majority of the questions having five ($525525$ items) or six distractors ($7171$ items). An example instance is provided in Table 1. As can be seen, questions of this dataset are longer ($755755$ characters compared to $103103$ characters for the Biopsychology set) and are of technical nature.
3.3 CMCQRD
The Cambridge Multiple-Choice Questions Reading Dataset (CMCQRD) consists of 120 text passages, each accompanied with a number of MCQs (total of 793 question items) aiming at evaluating the student’s language comprehension abilities. While the questions target various Common European Framework of Reference for Languages (CEFR) proficiency levels (B1 to C2), we do not leverage this additional information in our experiments. Furthermore, compared to the Biopsychology and USMLE datasets, question items in the CMCQRD are the longest, with an average of 3,61836183,6183 , 618 characters per item.
@@ 4 Approach
Given an MCQ, the task is to predict its difficulty measured by the proportion of students that select the correct choice (in addition to the IRT-based metric available for the CMCQRD dataset). An MCQ item consists of the stem/question, a single correct choice/answer and a number of incorrect choices/distractors (also known as “foils").
Figure 1 illustrates our approach to this task. Our design is centered around a simple Random Forest Regressor [^2] As provided by the Scikit-Learn Library, using the default hyper-parameters [18]. which receives as input a vectorised representation of the MCQ, as well the uncertainty of multiple LLMs answering the same MCQ [^3] Simple vector concatenation is used to combine the text and uncertainty features.. We opted for a relatively simple Random Forest Regressor, as it allows for analysis using explainability methods (through SHAP post-hoc analysis), while still effectively demonstrating the usefulness of model uncertainty in this context. As features, we use Textual Features and Model Uncertainty, as described in sections 4.1 and 4.2. We further supplement this basic system using higher-level textual features namely Linguistic Features and Choice Similarity, described in section 4.3.
@@ 4.1 Textual Features
TF-IDF Scores capture how important a word is to a document within a collection by balancing its frequency in that specific document against its rarity across all documents [27]. In the current context, we consider each question item (along with its choices) as a single document. To capture multi-word technical terms, such as “interstitial fibrosis", our analysis considers both individual words (unigrams) and two-word combinations (bigrams). Furthermore, we disregard terms that appear in more than 75% of documents, and only use the $10001000$ most important features (as determined by the TF-IDF values) to increase efficiency.
@@
Word embeddings are a technique whereby words are encoded as dense vectors in a continuous vector space, capturing semantic relationships between words. We evaluate two embedding approaches: General BERT embeddings [5] and domain-specific Bio-Clinical BERT embeddings [2]. The Bio-Clinical BERT embeddings, previously also employed by team ITEC in the BEA 2024 shared task [29], offer specialized medical domain text encoding that potentially encapsulates more accurately the semantic content of each question item. Both techniques yield a 768-dimensional vector representation. We only conduct experiments using Bio-Clinical BERT embeddings in the Biopsychology and USMLE datasets, as CMCQRD does not focus on the biology/medical domain.